OpenStudy (moongazer):

What is the difference between 5 4 11--- and 21--- ? 21 51

OpenStudy (moongazer):

give me the easiest solution please

OpenStudy (anonymous):

First one, keep 21 at bottom. Multiply 21 by 11; add answer to 5. Put that total on top.

OpenStudy (anonymous):

5 236 11----- = ------- 21 21 4 1075 21------ = ------- 51 51

OpenStudy (anonymous):

\[ 1171 \over 119\]

OpenStudy (anonymous):

\[21=3\times 7\] \[51=3\times 17\] \[\frac{5}{21}=\frac{5}{21}\times \frac{17}{17}=\frac{85}{357}\] \[\frac{4}{51}=\frac{4}{51}\times \frac{7}{7}=\frac{28}{357}\] \[\frac{85}{357}+\frac{28}{357}=\frac{113}{357}\] and \[11+21=32\] answer is \[32\tfrac{113}{357}\]

OpenStudy (anonymous):

Difference is 236 1075 ------ - ------- (LCM of 21 and 51 is 357) 21 51 4012 - 7525 - 3513 - 1171 -------------- = ---------- = --------- 357 357 119

OpenStudy (anonymous):

@satellite we have to find the difference and not the sum !!! :)

OpenStudy (moongazer):

the answer is 100 -9 --- 119 can this have an easier solution because this should be answered in just one minute with no calculators but it is multiple choice

OpenStudy (anonymous):

holy moly i added! sorry. it says "difference"

OpenStudy (moongazer):

i already know how to do it but i want an easier or shortcut but if it is not possible it is ok

OpenStudy (anonymous):

no shortcut at all because the answer is negative. gotta write out all the steps harkirat sent.

OpenStudy (moongazer):

ohh ok

OpenStudy (anonymous):

easier method cud b this just multiply the first fraction one by 51/51 and second fraction one by 21/21 then subtract the resulting fractions and reduce the answer to lowest form

OpenStudy (moongazer):

what?

OpenStudy (anonymous):

12036 - 22575 - 10539 - 1171 100 -------------- = --------- = ------- = - 9 ------- 21 x 51 1071 119 119

OpenStudy (anonymous):

clear now??? This will save the time taken to find the LCM....

OpenStudy (moongazer):

ohhh ok Maybe it is just a big number so i find it hard. I will just be practicing to be fast in doing this kind of fractions. I hope to be better in math now because i am loving it already and math is really an important subject. Thanks! :)

OpenStudy (anonymous):

Hi......The fractions are very small..If you need a practical answer then just find the difference between the integers(21-11=10) ;)

OpenStudy (moongazer):

what do you mean?

OpenStudy (anonymous):

for doing fractions quickly you should be thorough with finding LCM know your division rules for 2,3,4,5,6,7,8,9,10,11 etc.

OpenStudy (moongazer):

what?

OpenStudy (moongazer):

i am not that good at fractions because the teacher who taught me about fraction is hard to understand and i just understood it by my self as time goes by.

OpenStudy (anonymous):

5/21 and 4/51 are very small.....so you may neglect them...if you r answering an obejective type q it may help...

OpenStudy (anonymous):

@mmongazer practice as many sums as u can.....

OpenStudy (moongazer):

yup, practice is really the key to be good in math. :) Practice,practice and practice more as my teacher said. gtg i'll sleep already :)

OpenStudy (anonymous):

good night :))